# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

### 03 Dec 2023 v2

@@ 3.3 Neural Architecture Search (NAS) Framework for TransNAS-TSAD
The NAS framework shown in Figure 1 is pivotal in achieving our primary objective of effectively detecting anomalies in multivariate time series data. By employing the NSGA-II algorithm, our NAS process systematically navigates the vast architectural landscape of transformer models. This step is essential in identifying an architecture that not only excels in anomaly detection accuracy but also aligns with the computational constraints of practical deployment scenarios. The rigorous optimization and selection of transformer architectures through NAS directly contribute to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
@@ 3.3.1 Search Space Definition
The search space outlined in Table 1 within TransNAS-TSAD represents a diverse and extensive array of potential model architectures and hyperparameters. This assortment is meticulously designed to tailor neural network configurations specifically for time-series anomaly detection:
  • 1.
@@
Parameter Type Parameter Name Search Space Description
Training
Learning Rate
$1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale)
Rate at which the model learns.
Dropout Rate
0.1 to 0.5
Regularization to prevent overfitting.
Batch Size
16 to 128 (step of 16)
Number of samples per training step.
Gaussian Noise
$1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale)
Noise added for robustness.
Time Warping
True, False
Augmentation technique for time series.
Time Masking
True, False
Augmentation technique to mask intervals.
Window Size
10 to 30
Length of the input sequence window for the model.
Architectural
Positional Encoding Type
Sinusoidal, Fourier
Encoding type for sequence position.
Dimension Feedforward
8 to 128 (log scale)
Size of the feedforward network.
Encoder Layers
1 to 3
Number of layers in the encoder.
Decoder Layers
1 to 3
Number of layers in the decoder.
Activation Function
ReLU, Leaky ReLU, Sigmoid, Tanh
Non-linearity after each layer.
Attention Type
Scaled Dot Product
Type of attention mechanism.
Number of Attention Heads
Equal to feature dimension
Parallel attention layers.
Use Linear Embedding
True, False
Option to use a linear embedding layer.
Layer Normalization
Layer, Batch, Instance
Type of normalization used.
Self Conditioning
True, False
Conditioning strategy for the model.
Number of FFN Layers
1 to 3
Layers in the feedforward network.
Phase Type
1phase, 2phase, Iterative
Model’s reconstruction and refinement strategy.
@@ 3.3.2 TransNAS-TSAD Evaluation Strategy
Each trial within the TransNAS-TSAD NAS process entails a rigorous evaluation of models generated with varying architectural and training hyperparameters. This evaluation adopts a multi-objective approach, primarily focusing on two critical aspects: the F1 score and the number of model parameters. The F1 score serves as a key indicator of the model’s accuracy in anomaly detection, balancing precision and recall, while the number of parameters gauges the model’s architectural complexity and computational efficiency. The objective is to identify models that not only demonstrate high proficiency in accurately detecting anomalies (as reflected in a high F1 score) but also maintain a streamlined architecture (evidenced by a lower count of parameters). This dual-criteria assessment ensures the selection of models that are both effective in performance and practical in deployment, aligning with the overarching goal of achieving optimal anomaly detection with computational resourcefulness.
3.3.3 Multi-Objective Optimization with NSGA-II
TransNAS-TSAD employs the NSGA-II algorithm to balance the objectives of performance and computational efficiency. The optimization process is defined by the following equations:
$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})]$ (3)
@@ 3.3.3 Multi-Objective Optimization with NSGA-II
3.3.4 Pareto Front Exploration
@@
$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\}$ (6)
@@ 3.3.5 Utilization of the Pareto Front
Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II (Deb et al., (2002)) for the multi-objective optimization. This algorithm elucidates the sequential steps taken from initialization to final selection, highlighting the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
1 Initialization: Set up the Optuna framework with the NSGA-II sampler. Initialize the search with a population of transformer architectures from the predefined search space. Define evolutionary process parameters: population size, number of generations, mutation and crossover rates. Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time. for each study in Optuna do
2       for each trial in the study do
3             Define the transformer model architectural parameters and training hyperparameters within the search space constraints. Construct and train the transformer model on the Training dataset. Evaluate the model on the Validation dataset using the specified metrics. Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency. if early stopping criteria are met based on validation performance or predefined resource constraints then
4                   Terminate the trial prematurely.
5             end if
6
7       end for
8      Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions. Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
9 end for
Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
@@ 3.3.6 Evolutionary Process in TransNAS-TSAD
The evolutionary process within TransNAS-TSAD serves a critical role in aligning our methodology with the overarching goal of refining anomaly detection capabilities. Through iterative refinement and selection of model architectures, this process ensures that our models are not only accurate in identifying anomalies but also evolve to become more efficient and adaptable to various data characteristics. This continuous improvement and adaptation are key to achieving high-performance anomaly detection, which stands at the core of our research objectives.
  • 1.
    Generation Loop: New models are proposed, evaluated, and bred in each generation, with superior architectures refining the Pareto front.
  • 2.
@@ 3.4 Offline Optimization of Transformer Architecture for Anomaly Detection
    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of models from high-performance, resource-intensive ones to balanced and resource-efficient alternatives.
@@
3.4.1 Tailoring Architectures to Specific Deployment Needs
@@ 3.5 Advanced Anomaly Detection with Adversarial Elements in TransNAS-TSAD
TransNAS-TSAD represents a significant advancement in the field of anomaly detection in time series data, strategically incorporating and extending adversarial learning paradigms. Drawing inspiration from the two-phase approach of the TranAD model, our framework elevates this concept by introducing a neural architecture search (NAS) driven methodology. This approach dynamically selects between the refined two-phase approach and our novel iterative self-adversarial approach at each trial. The integration of these advanced techniques addresses the complexities inherent in time series data, particularly in capturing subtle and sophisticated anomalies that traditional encoder-decoder architectures might overlook.
Our development of the iterative self-adversarial approach, in particular, marks a notable enhancement over existing methods. By embedding elements of adversarial training within both the two-phase and iterative approaches, TransNAS-TSAD achieves a comprehensive and adaptable anomaly detection capability. This not only enhances the sensitivity and accuracy of the detection process but also ensures robustness and adaptability across diverse datasets. The synergy of these methodologies within TransNAS-TSAD aligns seamlessly with our research goal of developing a versatile and effective solution for time series anomaly detection, demonstrating our commitment to pushing the boundaries of current technologies.
3.5.1 Two-Phase Adversarial Approach
Drawing from adversarial learning paradigms, the two-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
@@ 3.5.1 Two-Phase Adversarial Approach
This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window. The preliminary input reconstruction yields a focus score, defined by the deviation of the reconstructed output from the actual input:
$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2}$
where $O_{\text{initial}}$ is the output from the first phase and $W$ is the input time-series window. The resultant “focus score,” derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
@@
Incorporating adversarial elements, this phase utilizes the focus scores from the phase 1 as a guiding beacon. In the adversarial phase, the second decoder aims to maximize the difference between its output and the input. Simultaneously, the first decoder aims to minimize this difference:
@@
3.5.2 Iterative Self-Adversarial Approach
Going beyond the structured two-phase reconstruction, TransNAS-TSAD introduces a dynamic iterative approach. This approach, embedded with self-adversarial mechanisms, continually refines its reconstructions.
3.5.3 Iterative Refinement:
@@
$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2}$
@@
$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|$
@@
3.5.4 Adaptable and Robust Anomaly Detection
@@ 3.6 Harnessing the Power of TransNAS-TSAD for Anomaly Detection
3.6.1 The Anomaly Scoring Mechanism
Every incoming data point, denoted as $\hat{W}$, is assessed for anomalies by computing a score that quantitatively reflects its deviation from expected patterns. In this context, $\hat{W}$ represents the new window of the time series data under consideration, and $\hat{C}$ embodies any associated contextual information that could influence the anomaly detection process. The anomaly score, $s$, is calculated as:
$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2}$ (7)
@@ 3.6.1 The Anomaly Scoring Mechanism
3.6.2 Dual Inference Pathways
@@
Figure 2 illustrates the anomaly detection process on dimension 0 of the SMD test dataset using TransNAS-TSAD. The upper plot displays the true data points (in blue) and the predicted values (in red), while the lower plot indicates the computed anomaly scores and their corresponding labels, offering a visual interpretation of the model’s detection capabilities in identifying anomalies.
@@
3.6.3 Dynamic Anomaly Thresholding: The Sentinel
@@
$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x)$ (8)
@@ 3.6.4 Augmentative Strategies for Enhanced Precision
$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}}$ (10)
@@ 4 Experimental Setup for TransNAS-TSAD
Our study rigorously evaluates the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD (Tuli et al., (2022)), LSTM-NDT (Hundman et al., (2018)), DAGMM (Zong et al., (2018)), OmniAnomaly (Su et al., (2019)), MSCRED (Zhang et al., (2019)), MAD-GAN (Li et al., (2019)), USAD (Audibert et al., (2020)), CAE-M (Zhang et al., (2021)), and GDN (Deng & Hooi, (2021)). To ensure a fair comparison, we utilize the hyperparameter configurations as specified in the original publications of these models, relying on their publicly available implementations.
The experimental infrastructure comprises a Google Colab Pro environment, leveraging a NVIDIA Tesla T4 GPU (16GB memory) and an Intel Xeon E5-2670 v3 CPU (8 cores, 51GB memory). The implementation is grounded in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling. These tools are instrumental in realizing the model implementation, executing the NSGA-II based NAS algorithm, and managing dataset operations.
A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design, navigating a comprehensive search space of architectural and training hyperparameters (detailed in Table 1). This search space is meticulously crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset.
Consistent with the practices in OmniAnomaly and TranAD, we apply the peaks over threshold (POT) method (Leadbetter, (1991)) across all datasets with a uniform coefficient of $10^{-4}$. The low quantile parameter, however, is fine-tuned for each specific dataset to maintain alignment with established benchmarks, thus ensuring an equitable basis for comparative analysis.

### 04 Mar 2024 v3

@@ 3.3 Neural Architecture Search (NAS) Framework for TransNAS-TSAD
The NAS framework shown in Figure 1 is pivotal in achieving our primary objective of effectively detecting anomalies in multivariate time series data. By employing the NSGA-II algorithm, our NAS process systematically navigates the vast architectural landscape of transformer models. This step is essential in identifying an architecture that not only excels in anomaly detection accuracy but also aligns with the computational constraints of practical deployment scenarios. The optimal selection of transformer architectures through NAS directly contributes to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
@@ 3.3.1 Search Space Definition
The search space within TransNAS-TSAD and outlined in Table 1 represents a diverse and extensive array of potential model architectures and hyperparameters. This assortment is meticulously tailored toward neural network configurations designed specifically for time-series anomaly detection:
@@
Parameter Type Parameter Name Search Space Description Training Learning Rate $1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale) Rate at which the model learns. Dropout Rate 0.1 to 0.5 Regularization to prevent overfitting. Batch Size 16 to 128 (step of 16) Number of samples per training step. Gaussian Noise $1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale) Noise added for robustness. Time Warping True, False Augmentation technique for time series. Time Masking True, False Augmentation technique to mask intervals. Window Size 10 to 30 Length of the input sequence window for the model. Architectural Positional Encoding Type Sinusoidal, Fourier Encoding type for sequence position. Dimension Feedforward 8 to 128 (log scale) Size of the feedforward network. Encoder Layers 1 to 3 Number of layers in the encoder. Decoder Layers 1 to 3 Number of layers in the decoder. Activation Function ReLU, Leaky ReLU, Sigmoid, Tanh Non-linearity after each layer. Attention Type Scaled Dot Product Type of attention mechanism. Number of Attention Heads Equal to feature dimension Parallel attention layers. Use Linear Embedding True, False Option to use a linear embedding layer. Layer Normalization Layer, Batch, Instance Type of normalization used. Self Conditioning True, False Conditioning strategy for the model. Number of FFN Layers 1 to 3 Layers in the feedforward network. Phase Type 1phase, 2phase, Iterative Model’s reconstruction and refinement strategy.
@@ 3.3.2 TransNAS-TSAD Evaluation Strategy
The evolutionary optimization strategy of TransNAS-TSAD entails a rigorous evaluation of models generated with varying architectural and training hyperparameters. The multi-objective evaluation focuses on two critical aspects: the F1 score and the number of model parameters. The F1 score serves as a key indicator of the model’s accuracy in anomaly detection, balancing precision and recall, while the number of parameters gauges the model’s architectural complexity and computational efficiency. The objective is to identify models that not only demonstrate high proficiency in accurately detecting anomalies (as reflected in a high F1 score) but also maintain a streamlined architecture (evidenced by a lower count of parameters). This dual-criteria assessment ensures the selection of models that are both effective in performance and practical in deployment, aligning with the overarching goal of achieving optimal anomaly detection with computational resourcefulness.
Multi-Objective Optimization with NSGA-II
TransNAS-TSAD employs the NSGA-II algorithm, a proven leader in balancing the objectives of performance and computational efficiency in multi-objective optimization. Our choice of NSGA-II is substantiated by recent benchmark studies, such as those presented in [45], which demonstrate its superiority in identifying optimal solutions across various search spaces. This evidence supports NSGA-II as an effective approach for optimizing the complex and competing demands of our model architecture. The optimization process in TransNAS-TSAD is governed by the following equations:
$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})],$ (3)
@@
Pareto Front Exploration and utilization
@@
$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\},$ (6)
@@
Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II [21] for the multi-objective optimization. This algorithm highlights, from initialization to final selection, the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
@@
1:  Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
2:  Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
3:  Initialization: Set up the Optuna framework with the NSGA-II sampler. Initialize the search with a population of transformer architectures from the predefined search space. Define evolutionary process parameters: population size, number of generations, mutation and crossover rates.
4:  Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time.
5:  for each study in Optuna do
6:     for each trial in the study do
7:        Define the transformer model architectural parameters and training hyperparameters within the search space constraints.
8:        Construct and train the transformer model on the Training dataset.
9:        Evaluate the model on the Validation dataset using the specified metrics.
10:        Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency.
11:        if early stopping criteria are met based on validation performance or predefined resource constraints then
12:           Terminate the trial prematurely.
13:        end if
14:     end for
15:     Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions.
16:     Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
17:  end for
18:  Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
@@ 3.3.3 Evolutionary Process in TransNAS-TSAD
The evolutionary process within TransNAS-TSAD is critical in achieving our overarching goal of refining anomaly detection capabilities. Through iterative refinement and selection of model architectures, this process ensures that our models are not only accurate in identifying anomalies but also evolve to become more efficient and adaptable to various data characteristics. This continuous improvement and adaptation are key to achieving high-performance anomaly detection, which stands at the core of our research objectives.
  • Generation Loop: New models are generated, evaluated, and passed on to each generation, with superior architectures refining the Pareto front.
@@ 3.4 Offline Optimization of Transformer Architecture for Anomaly Detection
    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of high-performance, resource-intensive models to more balanced and resource-efficient alternatives.
@@
Tailoring Architectures to Specific Deployment Needs
@@ 3.5 Advanced Anomaly Detection with Adversarial Elements in TransNAS-TSAD
TransNAS-TSAD represents a breakthrough in the field of time series anomaly detection by strategically enhancing adversarial learning paradigms. While it draws foundational inspiration from the TranAD model’s two-phase approach, our framework introduces: a neural architecture search (NAS) strategy that effectively incorporates a third, iterative self-adversarial phase. This tripartite approach enables dynamic selection among the conventional two-phase mechanism and our advanced iterative phase, optimizing detection capabilities for each specific trial.
The inception of this third phase signifies a substantial evolution in anomaly detection techniques. By embedding adversarial elements within all three phases, TransNAS-TSAD not only retains the strengths of the traditional encoder-decoder models but also introduces a level of adaptability and precision previously unattainable. This method excels in identifying subtle and complex anomalies, leveraging the iterative adversarial training to refine detection with each iteration.
3.5.1 Three-Phase Adversarial Approach
Drawing from adversarial learning paradigms, the three-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
@@ 3.5.1 Three-Phase Adversarial Approach
This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window, and yields a focus score. The latter is defined by the deviation of the reconstructed output from the actual input:
$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2},$
where $O_{\text{initial}}$ is the output from the first phase and $W$ is the input time-series window. The resultant “focus score”, derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
@@
Incorporating adversarial elements, this phase utilizes the focus scores from phase 1. In the adversarial phase, the second decoder aims to maximize the difference between its output and the input. Simultaneously, the first decoder aims to minimize this difference:
@@
Phase 3 -Iterative Self-Adversarial Approach
Going beyond the structured two-phase reconstruction inspired by the TranAD work, TransNAS-TSAD introduces a dynamic iterative approach. This approach, embedded with self-adversarial mechanisms, continually refines its reconstructions.
Iterative Refinement:
@@
$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2},$
@@
$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|.$
@@
Adaptable and Robust Anomaly Detection
@@ 3.6 Harnessing the Power of TransNAS-TSAD for Anomaly Detection
The Anomaly Scoring Mechanism
Every incoming data point, denoted as $\hat{W}$, is assessed for anomalies by computing a score that quantitatively reflects its deviation from expected patterns. Here, $\hat{W}$ represents the new window of time series data under consideration. The anomaly score, $s$, is calculated as:
$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2},$ (7)
@@
Dual Inference Pathways
@@
Figure 2 illustrates the anomaly detection process on dimension 0 of the Server Machine dataset (SMD) test dataset using TransNAS-TSAD. The upper plot displays the true data points (in blue) and the predicted values (in red), while the lower plot indicates the computed anomaly scores and their corresponding labels, offering a visual interpretation of the model’s detection capabilities in identifying anomalies.
@@
Dynamic Anomaly Thresholding: The Sentinel
@@
$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x),$ (8)
@@ 3.6.1 Augmentative Strategies for Enhanced Precision
$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}},$ (10)
@@ 4 Experimental Setup for TransNAS-TSAD
We evaluate the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD [59], LSTM-NDT [24], DAGMM [72], OmniAnomaly [61], MSCRED [73], MAD-GAN [48], USAD [4], CAE-M [74]), and GDN [19]. To ensure a fair comparison, we utilize the hyperparameter configurations as specified in the original publications of these models, relying on their publicly available implementations.
The experimental infrastructure comprises a Google Colab Pro environment, leveraging a NVIDIA Tesla T4 GPU (16GB memory) and an Intel Xeon E5-2670 v3 CPU (8 cores, 51GB memory). The NSGA-II based NAS algorithm and data management operations are implemented in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling.
A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design by navigating a comprehensive search space of architectural and training hyperparameters, as shown in Table 1. This search space is crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset. Consistent with the practices in OmniAnomaly and TranAD, we apply an enhanced version of the peaks over threshold (POT) method [49] across all datasets with a uniform coefficient of $10^{-4}$. This enhancement incorporates our previously described augmentative strategies, further fine-tuning the low quantile parameter for each specific dataset to align with established benchmarks and ensure equitable comparative analysis.